News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Saturday 9 October 1999 9.00am EST

During the last 24 hours Chandra exited the radiation belts, completed a post-belt CTI measurement and completed the GTO observation of EO102-72.3 with ACIS-S and the HETG. The GTO observation of AB Dor with ACIS-S and HETG was started at 7AM 10/9. The spacecraft systems continued to operate nominally.

Analysis of the ACIS CTI data from the radiation passage on day 275 shows that no further degradation in the energy resolution of the front-side illuminated CCD chips has occurred.

Return to Status Reports